One property Q of a material is used to describe the behaviour of sound waves in the material.Q is defined as the pressure P of the sound wave divided by the speed v of the wave and the surface area A of the material through which the wave travels:


Q=P/vA


1
Expert's answer
2021-09-28T16:55:03-0400

"\frac{kg \cdot m^{-2}}{m \cdot s^{-1} \cdot m^{2}} = kg \cdot s \cdot m^{-5}"